Mean life time of a sample of 400 of an product by a company is 1600 hours with a standard deviation of 150 hours. Test the hypothesis that the mean life of the product is more than 1570 hours at 1% level of significance


1
Expert's answer
2021-01-26T03:44:41-0500

The provided sample mean is "\bar{x}=1600" and the sample standard deviation is "s=150,"

and the sample size is "n=400."

The following null and alternative hypotheses need to be tested:

"H_0:\mu\leq1570"

"H_1:\mu>1570"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," the number of degrees of freedom are "df=n-1=400-1=399," and the critical value for a right-tailed test is "t_c=2.33573."

The rejection region for this right-tailed test is "R=\{t:|t|>2.33573\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{1600-1570}{150/\sqrt{400}}=4"

Since it is observed that "t=4>2.33573=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 1570, at the 1% significance level.

Using the P-value approach:

The p-value for "t=4, df=399, \alpha=0.01, right-tailed" is "p=0.000038," and since "p=0.00003<0.01=\alpha," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 1570, at the 1% significance level.